a. The solubility product, Ksp, for AgI is 8.3 x 10-17. What is the concentration of iodide ion (I- ) in solution saturated in silver iodide.

b. What would the concentration of silver(I) ion (Ag+ ) be for a saturate solution of AgI which is also 0.20 M in NaI (soluble, of course)
Answer #1

(a.) The equilibrium reaction in saturated solution of AgI is :

AgI (s) Ag+ (aq) + I- (aq)

Ksp AgI = [Ag+][I-]

Ksp AgI = (s) * (s)

where s is the molar solubility of AgI

8.3 x 10-17 = s2

s = (8.3 x 10-17)1/2

s = 9.1 x 10-9 M

Concentration of iodide, [I-] = s = 9.1 x 10-9 M

(b.) initial concentration of I- = [NaI] = 0.20 M

Ksp AgI = [Ag+][I-]

Ksp AgI = (s) * (s + 0.20 M)

where s is the molar solubility of AgI

8.3 x 10-17 = s2 + 0.20s

s2 + 0.20s - 8.3 x 10-17 = 0

Solving for s, s = 4.2 x 10-16 M

Concentration of Ag+ = [Ag+] = 4.2 x 10-16 M